OpenStudy (anonymous):

Inequality with variables in power: x^(x+6) < x^(4x-3) ------------------- same bases so: x+6 < 4x-3 9 < 3x 3 < x I got that answer right, but there is another answer that I don't know how to find which is "0

OpenStudy (anonymous):

\[x ^{\left( x+6 \right)}<x ^{\left( x+3 \right)}\] \[x ^{\left( x+3+3 \right)}-x ^{\left( x+3 \right)}<0,x ^{\left( x+3 \right)}\left( x^3-1 \right)<0\]

OpenStudy (anonymous):

both the factors are of different sign.

OpenStudy (anonymous):

if x>1,then both the factors are positive. if x=1 then the product is 0. so x<1

OpenStudy (anonymous):

\[hence x^3-1<0,so~ x ^{\left( x+3 \right)}>0\]

OpenStudy (anonymous):

Assuming the inequality was an equation, the answers would be x=3 and x=1, how would you find that x=1? Because 1 to any power is 1?

OpenStudy (anonymous):

if x>0,x<1.then \[x^3<1,x^3-1<0\]